# GL1-02-06-24-2-supplement

### v2

@@ 7.1 Estimation Without Misreporting
::=:= $\mathbb{I}\left[(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}),(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11})\in\Psi_{11}^{(1)};\;(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}),(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11})\in\Psi_{11}^{(0)}\right]$
×$\times\left(\left[\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(1)}}\left(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}\right)-\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(1)}}\left(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}\right)\right]-\left[\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}\left(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}\right)-\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}\left(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}\right)\right]\right),$
where $\widehat{\boldsymbol{\eta}}_{11}^{(0)}:=\widehat{\boldsymbol{\eta}}_{01}^{(0)}+\widehat{\boldsymbol{\eta}}_{10}^{(0)}-\widehat{\boldsymbol{\eta}}_{00}^{(0)}$ and $\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}(v):=\widehat{F}_{{\widetilde{\varepsilon}}_{10}^{(0)}}(v)$ for all $v\in\mathbb{R}$, while $\Psi_{11}^{(d)}$, $d\in\{0,1\}$, denote suitable trimming sets. Under regularity conditions on the different estimators, the above procedure yields consistent estimates of the DTTs of interest, see Section LABEL:sec:proofEstimation of the supplementary material for details. Finally, when the median (or mean) as well as the variance parameters actually differ, a slightly modified procedure can be adopted. In particular, we can construct estimators for the median (or mean) parameter $\eta_{c,gt}^{(d)}$ and the scaling parameter $\lambda_{gt}^{(d)}$ as follows:
@@
[…]
for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$, where $\Psi^{(d)}_{n,gt}$ now denotes a trimming set on which the index density is strictly bounded away from zero and that slowly expands with the sample size (cf. Hansen, 2008). This leads to an alternative estimator for the unscaled error term distribution:
$\widehat{F}_{{\varepsilon}_{gt}^{(d)}}(v):=\widehat{F}_{\widetilde{\varepsilon}_{gt}^{(d)}}\left(\widehat{\eta}_{c,gt}^{(d)}+\widehat{\lambda}_{gt}^{(d)}v\right),$
for all $\left(\widehat{\eta}_{c,gt}^{(d)}+\widehat{\lambda}_{gt}^{(d)}v\right)\in\Psi^{(d)}_{n,gt}$, though deriving the asymptotic distribution in this case is more involved and beyond the scope of the paper.
7.2 Estimation With Underreported Outcomes
In this subsection, we turn to the estimation of the DTTs when outcomes are subject to underreporting, as elaborated in Section 5. Acknowledging the non-constructive nature of the identification argument presented therein, our subsequent exposition assumes that $\overline{C}_{gt}^{(d)}(\cdot,\cdot)$ is a parametric copula, specifically $\overline{C}_{gt}^{(d)}\left(\cdot,\cdot;\rho_{gt}^{(d)}\right)$, parameterized by a dependence parameter $\rho_{gt}^{(d)}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$. Additionally, we postulate that the distributions of $-\varepsilon_{gt}^{(d)}$ and $-\nu_{gt}^{(d)}$ are known. For simplicity, we assume invariance across groups, time and treatment statuses in terms of copula functions and marginal distributions, i.e., $F_{\varepsilon_{gt}^{(d)}}=F_{\varepsilon}$, $F_{\nu_{gt}^{(d)}}=F_{\nu}$, and $\overline{C}_{gt}^{(d)}=\overline{C}$. This assumption can be relaxed straightforwardly if needed.
We define $\widetilde{\boldsymbol{\theta}}^{(d)}_{gt}:=\left(\overline{\boldsymbol{\eta}}^{(d)\prime}_{gt},\lambda^{(d)}_{gt},\overline{\boldsymbol{\pi}}^{(d)\prime}_{gt},\zeta^{(d)}_{gt},\rho^{(d)}_{gt}\right)^{\prime}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\},$ and aggregate these into $\widetilde{\boldsymbol{\theta}}:=\left(\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{00},\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{01},\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{10},\widetilde{\boldsymbol{\theta}}^{(1)\prime}_{11}\right)^{\prime}$. With this, the estimator in (7.1) is modified to:
$\widehat{\widetilde{\boldsymbol{\theta}}}:=$ $\arg\min_{\widetilde{\boldsymbol{\theta}}\in\mathcal{B}}\sum_{i=1}^{n}\sum_{j=0}^{2}\mathbb{I}\left[C_{i}=j\right]\left\{\sum_{(g,t)\neq(1,1)}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\times\right.$
$\left.\ln\left[\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{(0)}-\kappa_{j+1}}{\lambda_{gt}^{(0)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\pi}_{gt}^{(0)}-\iota_{j+1}}{\zeta_{gt}^{(0)}}\right);\rho_{gt}^{(0)}\right)-\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{(0)}-\kappa_{j}}{\lambda_{gt}^{(0)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\boldsymbol{\pi}}_{gt}^{(0)}-\iota_{j}}{\zeta_{gt}^{(0)}}\right);\rho_{gt}^{(0)}\right)\right]\right.$
$\left.+\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\ln\left[\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j+1}}{\lambda_{11}^{(1)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\pi}_{11}^{(1)}-\iota_{j+1}}{\zeta_{11}^{(1)}}\right);\rho_{11}^{(1)}\right)\right.\right.$ (23)
$\left.\left.-\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j}}{\lambda_{11}^{(1)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}-\iota_{j}}{\zeta_{11}^{(1)}}\right);\rho_{11}^{(1)}\right)\right]\right\}$
This estimator is conceptually similar to the first step of the estimation procedure described in Nguimkeu et al. (2019), which deals with the estimation of linear models with a partially observed, endogenous regressor within a group and a time period. Using the estimates, we construct the estimated conditional DTTs as outlined in Proposition 6:
$\widehat{\tau}^{c}_{11}(j\mid\mathbf{x})$ $:=\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$
$\widehat{\tau}^{r}_{11}(j\mid\mathbf{z})$ $:=\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(1)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j}}{\widehat{\zeta}^{(1)}_{11}}\right)\right]-\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(0)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j}}{\widehat{\zeta}^{(0)}_{11}}\right)\right]$
$\widehat{\overline{\boldsymbol{\eta}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\eta}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\eta}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\eta}}}_{00}^{(0)}}{\widehat{\lambda}_{00}^{(0)}/\widehat{\lambda}_{01}^{(0)}},\quad\widehat{\lambda}_{11}^{(0)}:=\frac{\widehat{\lambda}_{01}^{(0)}\widehat{\lambda}_{10}^{(0)}}{\widehat{\lambda}_{00}^{(0)}},\quad\widehat{\overline{\boldsymbol{\pi}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\pi}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\pi}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\pi}}}_{00}^{(0)}}{\widehat{\zeta}_{00}^{(0)}/\widehat{\zeta}_{01}^{(0)}}\quad\text{and}\quad\widehat{\zeta}_{11}^{(0)}:=\frac{\widehat{\zeta}_{01}^{(0)}\widehat{\zeta}_{10}^{(0)}}{\widehat{\zeta}_{00}^{(0)}}.$
We also explore the construction of “marginalized” DTTs for the treated group after treatment (see, e.g., Chernozhukov et al., 2013):
[…]
@@ 7.2 Estimation With Underreported Outcomes
Here, $\widehat{F}_{\mathbf{X}_{11}}(\cdot)$ represents an estimator of the covariate distribution. For instance:
@@
$\widehat{\tau}^{c}_{11}(j):=$ $\frac{1}{n_{11}}\sum_{i=1}^{n_{11}}\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\mathbb{I}\left[\mathbf{X}_{i}\in\mathcal{X}_{11}\right]\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right)$

### v3

@@ 4.2.2 Distributional Treatment Effects
$\widehat{\tau}^{r}(j\mid\mathbf{z})$ $:=\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(1)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j}}{\widehat{\zeta}^{(1)}_{11}}\right)\right]-\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(0)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j}}{\widehat{\zeta}^{(0)}_{11}}\right)\right]$
$\widehat{\overline{\boldsymbol{\eta}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\eta}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\eta}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\eta}}}_{00}^{(0)}}{\widehat{\lambda}_{00}^{(0)}/\widehat{\lambda}_{01}^{(0)}},\quad\widehat{\lambda}_{11}^{(0)}:=\frac{\widehat{\lambda}_{01}^{(0)}\widehat{\lambda}_{10}^{(0)}}{\widehat{\lambda}_{00}^{(0)}},\quad\widehat{\overline{\boldsymbol{\pi}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\pi}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\pi}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\pi}}}_{00}^{(0)}}{\widehat{\zeta}_{00}^{(0)}/\widehat{\zeta}_{01}^{(0)}}\quad\text{and}\quad\widehat{\zeta}_{11}^{(0)}:=\frac{\widehat{\zeta}_{01}^{(0)}\widehat{\zeta}_{10}^{(0)}}{\widehat{\zeta}_{00}^{(0)}}$
are the corresponding ML estimates. Similarly, the construction of “marginalized” DTTs for the treated group after treatment (see, e.g., Chernozhukov et al., 2013) are:
@@
Here, $\widehat{F}_{\mathbf{X}_{11}}(\cdot)$ represents an estimator of the covariate distribution.
Remark 2: While Proposition 2 does not allow any conclusions about the relationship of $\overline{\boldsymbol{\eta}}_{gt}$ and $\lambda_{gt}$ across pre-treatment periods, note that the ML estimator mentioned in previous remark can be used to conduct “pre-treatment” tests of the CiC model structure in analogy to the popular “pre-trend” test. That is, using the consumption equation as illustration, we may test the implied CiC relationship in Equation (11) from Proposition 2 in pre-treatment periods $t=0,-1,-2,\ldots$. Specifically, under the assumption that Equation (11) holds also for any given pre-treatment period $t\in\{0,-1,-2,\ldots\}$, we may test:
@@
$H_{0}:\overline{\boldsymbol{\eta}}^{(0)}_{1t}-\overline{\boldsymbol{\eta}}^{(0)}_{0t}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}\left(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)}\right)=\mathbf{0}\quad\text{and}\quad\frac{\lambda^{(0)}_{1t}}{\lambda^{(0)}_{1(t-1)}}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}=0,$
against